# Decoding-Time Language Model Alignment with Multiple Objectives

### v1

@@ Abstract
Aligning language models (LMs) to human preferences has emerged as a critical pursuit, enabling these models to better serve diverse user needs. Existing methods primarily focus on optimizing LMs for a single reward function, limiting their adaptability to varied objectives. Here, we propose multi-objective decoding (MOD), a decoding-time algorithm that outputs the next token from a linear combination of predictions of all base models, for any given weightings over different objectives. We exploit a common form among a family of $f$-divergence regularized alignment approaches (such as PPO, DPO, and their variants) to identify a closed-form solution by Legendre transform, and derive an efficient decoding strategy. Theoretically, we show why existing approaches can be sub-optimal even in natural settings and obtain optimality guarantees for our method. Empirical results demonstrate the effectiveness of the algorithm. For example, compared to a parameter-merging baseline, MOD achieves 12.8% overall reward improvement when equally optimizing towards $3$ objectives. Moreover, we experiment with MOD on combining three fully-finetuned LLMs of different model sizes, each aimed at different objectives such as safety, coding, and general user preference. Unlike traditional methods that require careful curation of a mixture of datasets to achieve comprehensive improvement, we can quickly experiment with preference weightings using MOD to find the best combination of models. Our best combination reduces toxicity on Toxigen to nearly 0% and achieves 7.9–33.3% improvement across other three metrics (i.e., Codex@1, GSM-COT, BBH-COT).
@@ 1 Introduction
Figure 1: Multi-objective decoding. We prepare LMs tuned for each objective in advance. Then given preference weightings $w$, input prompt $x$ and context $y_{<t}$, $y_{t}$ is greedily decoded from an algebraic combination of predicted probabilities from each LM, achieving precise control over each objective.
@@
In this work, we tackle the question: Given a set of policies corresponding to different rewards and linear coefficients for the rewards, can we find a training-free policy corresponding to the interpolated reward? We introduce multi-objective decoding (MOD; see Figure 1), which combines the predictive distributions of individual models trained for single objectives. This approach is inspired by Legendre transform in convex optimization convexoptimization , which allows us to derive a closed-form solution from a family of $f$-divergence regularized optimization approaches RLHF ; DPO ; fDPO (e.g., PPO, DPO are optimizing for the reward function with KL-divergence penalty), and its efficient approximation. The resulting method extends prior work employing logit arithmetic for decoding-time alignment proxytuning ; Jailbreak ; huang2024offset ; liu2024decoding , but we are the first to successfully achieve decoding towards multiple objectives simultaneously. We compare the design of our approach with existing multi-objective alignment baselines in Table 1.
Table 1: Overall comparison with other baselines. “Free from RM” refers to not requiring reward models. “Free from prompting” refers to not requiring preference-driven prompts during inference. Generally, the number of preferences is much larger than the number of objectives here. Among them, our approach is the most versatile solution, with an acceptable cost.
@@
Importantly, our approach allows users to achieve arbitrary weightings of objectives at inference time, avoiding the need for extensive retraining iterations. Additionally, our approach offers users more precise and interpretable control over the customization of AI outputs, thereby enhancing both personalization and performance. We conduct experiments across various tasks including Reddit Summary nisan2020summarize , Helpful Assistant Bai2022TrainingAH , and Safety Alignment beavertails . Notably, our method can combine models of different scales, and it is effective not only for RLHF and DPO models but also can be extended to supervised fine-tuned (SFT) models. This insight is supported by experiments on combining 13B DPO models and a 7B SFT model for Open Instruction-Following wang2023far ; ivison2023camels task.
@@
  • We introduce a training-free, simple yet effective algorithm, MOD, for multi-objective alignment of language models. Given strong-barrier function regularized base policies trained for a single objective, we are able to derive and efficiently decode a closed-form solution for an interpolated objective with optimality guarantees, based on Legendre transformation. Notably, our comprehensive framework generalizes and explains many existing tuning approaches and decoding strategies proxytuning ; Jailbreak ; huang2024offset ; liu2024decoding ; modpo . See Section 4.
  • In extensive experiments, we demonstrate the strong performance of MOD. For instance, compared to parameter merging, MOD achieves a $12.8$% overall reward improvement when equally optimizing towards three objectives on Helpful assistant task. When combining $3$ Tülu models, our best configuration significantly reduces Toxigen to nearly zero and results in a $7.9$% to $33.3$% improvement across the other three metrics (Codex@1, GSM-COT, BBH-COT). Additionally, experiments validate that our framework is applicable to SFT models and is still effective given a mix of positive and negative weightings (a case where the traditional training-free baseline does not work), showing its steerability. See Section 5.
@@
2 Related Work
Algorithms for aligning LMs to human preferences. The standard RLHF (PPO) approach InstructGPT ; nisan2020summarize ; wu2023fine optimizes over rewards with Reverse KL-divergence as divergence penalty, where the reward models are learned from human preference datasets. DPO DPO leverages the Bradley-Terry assumption BTmodel to directly optimize the same objective on preferences, in a supervised manner. $\Psi$-PO IPO further modifies the reward term to be optimized as other mappings from preference pairs; f-DPO fDPO replaces Reverse KL-divergence with other divergence measures. In addition, there are other efforts exploring alternative objectives and frameworks: SLiC-HF SLiC ; SLiCHF refer to the alignment process as sequence likelihood calibration; SPIN SPIN iteratively improves the model by leveraging synthetically generated data, thereby circumventing the need for human feedback; OPO alignonthefly employs established norms as constraints, achieving training-free alignment; and Lyu et  al. Lyu2024KeepingLA highlight the crucial role of prompt templates. In this work, we mainly focus on RLHF (PPO), DPO and their extensions.
Decoding-time algorithms for controllable generation. Response-level decoding algorithms sample a whole output $y$ from an anticipated probability distribution $p$. To solve this, energy-based methods are adopted in many works cold ; gradientbasedsampling , which involves continuous optimization for LMs to obtain gradient information. Besides, it can be viewed as maximizing $\log p(y)$ while satisfying some constraints, and Kumar et  al. Kumar2021ControlledTG utilizes simultaneous gradient descent to solve the dual problem. Token-level decoding algorithms decode token $y_{t}$ at timestep $t$, and are usually more efficient. Among them, Mudgal et  al. controlleddecoding , Liu et  al. Liu2023DontTA deploy value models to guide the decoding process; DeRa liu2024decoding works on hyper-parameter re-alignment and proposes the potential of a special case of MOD, while introducing a per-token distribution approximation; proxy-tuning proxytuning ; Jailbreak ; huang2024offset tunes a small model and applies it to steer a larger base model by operating on logits.
Multi-objective LMs alignment. Multi-objective alignment is the task of aligning language models to multiple objectives simultaneously. This is important for mitigating the dichotomy between different dimensions Vamplew2017HumanalignedAI ; Bai2022TrainingAH and catering to the diverse needs of users personalizedsoup ; dong-etal-2023-steerlm . Approaches for multi-objective alignment fall into the following categories: 1) Retraining. The most natural approach to solve multi-objective alignment is to retrain for a linearly combined multiple reward functions (MORLHF wu2023fine ; Bai2022TrainingAH ). And MODPO modpo enables the model to align with multi-objective on the initial preference dataset, by integrating a learned reward representation. 2) Parameter-merging. A line of work rewardedsoups ; personalizedsoup ; Lin2023MitigatingTA , represented by rewarded soups (RS), establishes a paradigm aimed at providing a training-free solution which obtains weights of the policy as a linear combination of weights of trained policies for each single objective, inspired by modelsoup and its other applications WARM ; lawson2023merging . 3) Preference-conditioned prompting. The preference-conditioned learning approaches Zhu2023ScalingPD ; Basaklar2022PDMORLPM train a policy conditioned on preference weightings to maximize the expected rewards, and are reflected in LMs alignment as preference-conditioned prompting: this line of work RiC ; haoxiang2024arithmetic ; guo2024controllable directly present the preference weightings in prompts after a fine-tuning process. The latter two paradigms are more efficient, while relying heavily on either reduced mis-specification hypothesis rewardedsoups or unguaranteed OOD generalization ability Zhou2023FreeFB , posing challenges in terms of interpretability and robustness.
@@ 3 Preliminaries
There are various ways of defining “multi-objective”. In this paper, we take a multi-objective reward function perspective. In this section, we will first give a formal definition of multi-objective reward functions. After that, because we focus exclusively on decoding by combining the predictions of a set of existing single-objective aligned LMs, we will give a formal assumption on each base LM considered in this paper. Finally, we will show the mathematical advantage of those base LMs under such assumptions. See full notations in Appendix B.
@@
Single objective alignment with $f$-divergence regularization. Each policy $\pi_{i}$ has been optimized for the corresponding reward function $\mathcal{R}_{i}$. However, it is well known that greedily optimizing towards maximum rewards can lead to over-optimization and worsen model performance rewardoveroptimization . Therefore, regularization has been incorporated to avoid large deviations from the reference policy. Alignment with KL-divergence regularization has been established as a standard formulation InstructGPT ; nisan2020summarize ; wu2023fine ; DPO ; xiong2024iterative ; theorynlhf . Recently, a sequential line of work fDPO ; tang2024fine has proposed replacing Reverse KL-divergence with a set of $f$-divergences such as Forward KL-divergence, JSD, and $\alpha$-divergence, which they claim can enhance generation diversity and decrease the expected calibration error guo2017calibration empirically. We observe that all these methods can be analyzed under the framework of $f$-divergences, where $f$ is a barrier function (see 1 and 2 in appendix for formal definitions). The closed form of each single-objective aligned LM $\pi_{i}$ can be written as:
@@
Strong-barrier function benefits multi-objective decoding. As discussed above, existing works choose different $f$ primarily to achieve different regularization behaviors. However, there is an extra benefit in the decoding setting: if the barrier function $f$ is continuously differentiable and strongly convex on $\mathbb{R}_{+}$, we can obtain a closed-form bijection between any single-objective aligned LM $\pi_{i}$ and the corresponding $\mathcal{R}_{i}$ as shown below (initially proposed in fDPO , see detailed proof in 1):
@@
where $Z_{i}(x)$ is the normalization factor with respect to $x$. In other words, rewards $\mathcal{R}_{i}$ are partially reversible when only $\pi_{i}$ are given. Crucially, such closed forms directly result in a possible linear combination of different outputs of $\{\pi_{i}\}_{i=1}^{M}$, as we will show in our main algorithm. In the rest of the paper, we call an $f$ with such properties a strong-barrier function.
Formal problem formulation. Given all those preliminaries, now we are ready to state our formal problem formulation: We are given a reference policy $\pi_{\operatorname{ref}}$ and a set of base policies $\{\pi_{i}\}_{i=1}^{M}$ trained for reward functions $\{\mathcal{R}_{i}\}_{i=1}^{M}$ under $f$-divergence regularization. On the other hand, we are unable to access $\mathcal{R}_{i}$ directly. Can we find a decoding algorithm such that, for any given preference weightings $w\in\Delta^{M-1}$ and input $x$, we can obtain a optimal response $y$ for the weighted multi-objective reward function $r(y|x)=\sum_{i=1}^{M}w_{i}\cdot\mathcal{R}_{i}(y|x)$, that is regularized by the reference policy?
@@ 4.1 Warm-up: an inefficient decoding version
where $C_{1}\in\mathbb{R}_{+}$ is some threshold constant. Now by leveraging the bijection property from strong-barrier function, as shown in Eq. (3), there exists a naive decoding format $\pi^{\star}$ for the dual problem (see detailed proof in 1):
@@
where $Z(x)$ and $Z^{\star}(x)$ are normalization factors. With this format, we can directly combine the outputs from $\{\pi_{i}\}_{i=1}^{M}$ during decoding. Unfortunately, computing the exact value of the normalization factor is nearly impossible as it requires looping over all possible $y$ in the output space.
@@ 4.2 Towards an efficient algorithm: reformulation and approximation
where $C_{2}\in\mathbb{R}_{+}$ is some threshold constant. Based on this observation and Legendre transform in convex optimization convexoptimization , we prove our key theorem which successfully gets rid of normalization factor and leads to the MOD algorithm, as follows (see detailed proof in Appendix D.3).
@@ 4.3 Main algorithm: efficient decoding with optimality for strong-barrier function
$y_{t}=\underset{s\in\Sigma}{\operatorname*{argmax}}\;\pi_{\operatorname{ref}}(y_{<t},s|x)\cdot(\nabla f)^{(-1)}\left(\sum_{i=1}^{M}w_{i}\cdot\nabla f\left(\frac{\pi_{i}(y_{<t},s|x)}{\pi_{\operatorname{ref}}(y_{<t},s|x)}\right)\right)\;.$ (6)
@@
$y_{t}=\operatorname*{argmax}_{s\in\Sigma}\;\prod_{i=1}^{M}\pi_{i}^{w_{i}}(y_{<t},s|x)\;,$ (7)
@@
Comparisons with other baselines. Our algorithm is significantly more efficient than retraining-based algorithms. In practice, the number of objectives is easily enumerable (e.g., $<5$ in Wang2023HelpSteerMH ; ultrafeedback ), making it feasible to finetune an LLM for each objective. In contrast, the number of preferences cannot be bounded due to the variability among users Casper2023OpenPA , which suggests that retraining-based algorithms like MORLHF and MODPO need to compute an impractical amount of times in order to match the preference of every user. Regarding memory efficiency, MOD requires loading multiple models simultaneously, which consume relatively higher memory cost. However, we mitigate this cost by ensembling a set of low-rank adapters or using distributed deployment in implementation. A comprehensive comparison with these baselines is shown in Table 1.
@@ 5.1 Experiment setup
Baselines. Rewarded soups (RS) rewardedsoups linearly merges each model’s parameters according to preference weightings, as $\theta=\sum_{i=1}^{M}w_{i}\cdot\theta_{i}$, where $\theta_{i}$ denotes the parameters of $\pi_{i}$. MORLHF wu2023fine optimizes for the weighted multi-objective reward function $\sum_{i=1}^{M}w_{i}\cdot\mathcal{R}_{i}$ using PPO, with the same configurations as training for single objective. MODPO modpo uses $\pi_{1}$’s output as an implicit reward signal of $\mathcal{R}_{1}$ and inserts it into the DPO objective for $\mathcal{R}_{2}$ to optimize for $w_{1}\mathcal{R}_{1}+w_{2}\mathcal{R}_{2}$, with the same configurations as training for single objective.

### v3

@@ Abstract
Aligning language models (LMs) to human preferences has emerged as a critical pursuit, enabling these models to better serve diverse user needs. Existing methods primarily focus on optimizing LMs for a single reward function, limiting their adaptability to varied objectives. Here, we propose multi-objective decoding (MOD), a decoding-time algorithm that outputs the next token from a linear combination of predictions of all base models, for any given weighting over different objectives. We exploit a common form among a family of $f$-divergence regularized alignment approaches (such as PPO, DPO, and their variants) to identify a closed-form solution by Legendre transform, and derive an efficient decoding strategy. Theoretically, we show why existing approaches can be sub-optimal even in natural settings and obtain optimality guarantees for our method. Empirical results demonstrate the effectiveness of the algorithm. For example, compared to a parameter-merging baseline, MOD achieves 12.8% overall reward improvement when equally optimizing towards $3$ objectives. Moreover, we experiment with MOD on combining three fully-finetuned LMs of different model sizes, each aimed at different objectives such as safety, coding, and general user preference. Unlike traditional methods that require careful curation of a mixture of datasets to achieve comprehensive improvement, we can quickly experiment with preference weightings using MOD to find the best combination of models. Our best combination reduces toxicity on Toxigen to nearly 0% and achieves 7.9–33.3% improvement across three other metrics (i.e., Codex@1, GSM-COT, BBH-COT).
@@ 1 Introduction
Figure 1: Multi-objective decoding. We prepare LMs tuned for each objective in advance. Then, given preference weightings $w$, input prompt $x$ and context $y_{<t}$, $y_{t}$ is greedily decoded from an algebraic combination of predicted probabilities from each LM, achieving precise control.
@@
In this work, we tackle the question: Given a set of policies corresponding to different rewards and linear coefficients for the rewards, can we find a training-free policy corresponding to the interpolated reward? We introduce multi-objective decoding (MOD; see Figure 1), which combines the predictive distributions of individual models trained for single objectives. This approach is inspired by Legendre transform in convex optimization convexoptimization , which allows us to derive a closed-form solution from a family of $f$-divergence regularized optimization approaches RLHF ; DPO ; fDPO (e.g., PPO, DPO are optimizing for the reward function with KL-divergence penalty), and its efficient approximation. The resulting method extends prior work employing logit arithmetic for decoding-time alignment proxytuning ; Jailbreak ; huang2024offset ; liu2024decoding , but we are the first to successfully achieve decoding towards multiple objectives simultaneously. We compare the design of our approach with existing multi-objective alignment approaches in Table 1.
Table 1: Overall comparison with other approaches. “Free from RM” refers to not requiring reward models. “Free from prompting” refers to not requiring preference-driven prompts during inference. Generally, the number of preferences is much larger than the number of objectives here. Among them, our approach is the most versatile solution.
@@
Importantly, our approach allows users to achieve arbitrary weightings of objectives at inference time, avoiding the need for extensive retraining iterations. Additionally, our approach offers users more precise and interpretable control over the customization of AI outputs, thereby enhancing both personalization and performance. We conduct experiments across various tasks including Reddit Summary nisan2020summarize , Helpful Assistant Bai2022TrainingAH , and Safety Alignment beavertails . Notably, our method can combine models of different scales, and it is effective not only for PPO and DPO models but also can be extended to supervised finetuned (SFT) models. This insight is supported by experiments on combining 13B DPO models and a 7B SFT model for Open Instruction-Following wang2023far ; ivison2023camels .
@@
  • We introduce a training-free, simple, yet effective algorithm, MOD, for multi-objective alignment of language models. Given strong-barrier function regularized base policies trained for a single objective, we are able to derive and efficiently decode a closed-form solution for an interpolated objective with optimality guarantees, based on Legendre transformation. Notably, our comprehensive framework generalizes and explains many existing tuning approaches and decoding strategies proxytuning ; Jailbreak ; huang2024offset ; liu2024decoding ; modpo . See Section 3.
  • In extensive experiments, we demonstrate the strong performance of MOD. For instance, compared to parameter merging, MOD achieves a $12.8$% overall relative reward improvement when equally optimizing towards three objectives on Helpful assistant task. When combining $3$ Tülu models, our best configuration significantly reduces Toxigen to nearly zero and results in a $7.9$% to $33.3$% relative improvement across the other three metrics (Codex@1, GSM-COT, BBH-COT). Additionally, experiments validate that our framework is applicable to SFT models and is still effective for given a mix of positive and negative weights (a case where the traditional training-free baseline does not work), showing its steerability. See Section 4.
@@ 2 Preliminaries
There are various ways of defining “multi-objective.” In this paper, we take a multi-objective reward function perspective. In this section, we will first give a formal definition of multi-objective reward functions. After that, because we focus exclusively on decoding by combining the predictions of a set of existing single-objective aligned LMs, we will give a formal assumption on each base LM considered in this paper. Finally, we will show the mathematical advantage of those base LMs under such assumptions. Notation is given in Appendix B.
@@
Single objective alignment with $f$-divergence regularization. Each policy $\pi_{i}$ has been optimized for the corresponding reward function $\mathcal{R}_{i}$. However, it is well known that greedily optimizing towards maximum rewards can lead to over-optimization and worsen model performance rewardoveroptimization . Therefore, regularization has been incorporated to avoid large deviations from the reference policy. Alignment with KL-divergence regularization has been established as a standard formulation InstructGPT ; nisan2020summarize ; wu2023fine ; DPO ; xiong2024iterative ; theorynlhf . Recently, a sequential line of work fDPO ; tang2024fine has proposed replacing Reverse KL-divergence with a set of $f$-divergences such as Forward KL-divergence, JSD, and $\alpha$-divergence, which they claim can enhance generation diversity and decrease the expected calibration error guo2017calibration empirically. We observe that all these methods can be analyzed under the framework of $f$-divergences, where $f$ is a barrier function (see 1 and 2 in Appendix D.1 for formal definitions). The closed form of each single-objective aligned LM $\pi_{i}$ can be written as:
@@
Strong-barrier function benefits multi-objective decoding. As discussed above, existing works choose different $f$ primarily to achieve different regularization behaviors. However, there is an extra property: if the barrier function $f$ is continuously differentiable and strongly convex on $\mathbb{R}_{+}$, we can obtain a closed-form bijection between any single-objective aligned LM $\pi_{i}$ and the corresponding $\mathcal{R}_{i}$ as shown below (initially proposed in fDPO , see detailed proof in 1):
@@
where $Z_{i}(x)$ is the normalization factor with respect to $x$. In other words, given the rewards and a prompt $x$, there is a closed form for the optimal policy, and given the optimal policies and $x$, there is a closed form for the rewards for every $y$. Crucially, such closed forms directly result in a possible linear combination of different outputs of $\{\pi_{i}\}_{i=1}^{M}$, as we will show in our main algorithm. In the rest of the paper, we call an $f$ with such properties a strong-barrier function.
Formal problem formulation. Given all those preliminaries, now we are ready to state our formal problem formulation: We are given a reference policy $\pi_{\operatorname{ref}}$ and a set of base policies $\{\pi_{i}\}_{i=1}^{M}$ trained for reward functions $\{\mathcal{R}_{i}\}_{i=1}^{M}$ under $f$-divergence regularization. And we assume that we are unable to access $\mathcal{R}_{i}$ directly. Can we find a retraining-free decoding algorithm such that, for any given preference weightings $w\in\Delta^{M-1}$ and input $x$, we can obtain an optimal response $y$ for the weighted multi-objective reward function $r(y|x)=\sum_{i=1}^{M}w_{i}\cdot\mathcal{R}_{i}(y|x)$, that is regularized by $\pi_{\operatorname{ref}}$?
@@ 3.1 Warm-up: an inefficient decoding version
where $C_{1}\in\mathbb{R}_{+}$ is some threshold constant. Now by leveraging the bijection property of a strong-barrier function, as shown in Eq. (3), there exists a naive decoding format $\pi^{\star}$ for the dual problem (see detailed proof in 1):
@@
where $Z(x)$ and $Z^{\star}(x)$ are normalization factors. With this form, we can directly combine the outputs from $\{\pi_{i}\}_{i=1}^{M}$ during decoding. Unfortunately, computing the exact value of the normalization factor is nearly impossible as it requires looping over all possible $y$ in the output space.
@@ 3.2 Towards an efficient algorithm: reformulation and approximation
where $C_{2}\in\mathbb{R}_{+}$ is some threshold constant. Based on this observation and Legendre transform in convex optimization convexoptimization , we prove our key theorem which gets rid of the normalization factor and leads to the MOD algorithm, as follows (see detailed proof in Appendix D.3).
@@ 3.3 Main algorithm: efficient decoding with optimality for strong-barrier function
$\underset{s\in\Sigma}{\operatorname*{argmax}}\;\pi_{\operatorname{ref}}(y_{<t},s|x)\cdot(\nabla f)^{(-1)}\left(\sum_{i=1}^{M}w_{i}\cdot\nabla f\left(\frac{\pi_{i}(y_{<t},s|x)}{\pi_{\operatorname{ref}}(y_{<t},s|x)}\right)\right)\;.$ (6)
@@
$\operatorname*{argmax}_{s\in\Sigma}\;\prod_{i=1}^{M}\pi_{i}^{w_{i}}(y_{<t},s|x)\;,$ (7)
@@
Comparisons with other approaches. Our algorithm is significantly more efficient than retraining-based algorithms. In practice, the number of objectives is easily enumerable (e.g., $<5$ in Wang2023HelpSteerMH ; ultrafeedback ), making it feasible to finetune an LM for each objective. In contrast, the number of preferences cannot be bounded due to the variability among users Casper2023OpenPA , which suggests that retraining-based algorithms like MORLHF and MODPO need to compute an impractical amount of times in order to match the preference of every user. Regarding memory efficiency, MOD requires loading multiple models simultaneously, which consume relatively higher memory cost. However, we mitigate this cost by ensembling a set of low-rank adapters or using distributed deployment in implementation. A comprehensive comparison with these baselines is shown in Table 1.
@@ 4.1 Experiment setup
Baselines. We adopt the representative parameter-merging method and retraining approaches as our baselines. Rewarded soups (RS) rewardedsoups linearly merge each model’s parameters according to preference weightings, as $\theta=\sum_{i=1}^{M}w_{i}\cdot\theta_{i}$, where $\theta_{i}$ denotes the parameters of $\pi_{i}$. MORLHF wu2023fine optimizes for the weighted multi-objective reward function $\sum_{i=1}^{M}w_{i}\cdot\mathcal{R}_{i}$ using PPO, with the same configurations as training for single objective. MODPO modpo uses $\pi_{1}$’s output as an implicit reward signal of $\mathcal{R}_{1}$ and inserts it into the DPO objective for $\mathcal{R}_{2}$ to optimize for $w_{1}\mathcal{R}_{1}+w_{2}\mathcal{R}_{2}$, with the same configurations as training for single objective.